# A Little Human Data Goes A Long Way

### v1

@@ 9 Ethical Considerations
The usage of synthetic data has several important ethical considerations. In the era of LMs trained on internet-wide corpora having poor documentation as to their exact data sources, it becomes challenging to ensure the privacy of individuals whose data may be obtainable via a public crawl (Yao et al., 2024). Additionally, models trained on massive internet-based data sources may contain implicit biases, illegal and/or highly offensive material that is hard to audit and clean (Bender et al., 2021). This data affects the synthetic data obtained from prompt models, and could unknowingly impose cultural or ethical viewpoints that are unintended or not well aligned with the use case in mind. Specifically, prior work has shown that one of the prompt models studied in this work, GPT-3.5, often disagrees with humans on key ethical questions (Felkner et al., 2024). The endeavour to completely replace human annotation with synthetic data generation also has key implications on the extent to which the field of NLP employs human annotators. It is possible that an increasing reliance on purely synthetic data reduces the demand for human annotation, which would place a downward pressure on the working standards and compensation awarded to the remaining human annotators (Weidinger et al., 2022). We argue in this work that we should not try to eliminate human annotation from our dataset and method design, showing that their work contributes uniquely helpful data points.
@@ 10 Acknowledgements
This work was funded by the Defense Advanced Research Projects Agency with award HR00112220046.
@@ Appendix A Supplemental Figures
We present a more detailed set of figures and tables to supplement the results presented in the main text.
@@
Synthetic Replacement Experiments: For figures in the main text where only one task is shown (Figure 1 and Figure 2), we provide the complete figures with both tasks (Figure 5 and Figure 6). We also provide the individual performance curves for these experiments (Figure 7 and Figure 8).
@@
To verify the robustness of the results, we show that the QA results are not an artifact of the choice of metric (Figure 1 and Figure 2), and that the same trends can be seen (Figure 9, Figure 10 and Figure 11) when using a different fine-tuning model (Mistral-7B), a more powerful prompting model (GPT-4-Turbo) and a more sophisticated prompting strategy (Chain-of-Thought Prompting). Across all configurations, we see a consistent decrease in performance when moving from 95% to 100% synthetic data, confirming that models trained on purely synthetic data can be improved by including just 125 real data points. For Chain-of-Thought Prompting, the authors manually annotated 3 examples with rationales per dataset to serve as the prompts. The complete examples and pipeline will be provided with the code once the review period is concluded.
@@
We additionally show that these trends hold across data scales (Figure 12 and Figure 13), replicating the experiment with n=3000 and n=1000. While the trend is clearly visible in both cases, the results for n=1000 have more variance and hence have a minority of cases where the relationship does not hold.
@@
Tradeoff Experiment: The main text shows results for the experiment detailed in Section 4 on the WANLI dataset (Figure 3), here we show results on the remaining three datasets (Figure 14) and provide (Figure 3)the number of additional synthetic points needed to match the performance gains of 200 additional real points (average, median and standard deviation for each dataset). ROPES shows similar results to WANLI, however FairyTaleQA and FEVER present different trends. On FEVER, we are able to reach the saturation point, after which additional data (whether synthetic or real) does not increase performance. Even in this case, we are able to reach this point of diminishing marginal return more rapidly when using a small amount of synthetic data. On a base synthetic training set of size $3000$, adding $200$ real data points drives the test accuracy to $89.25\%$, a score that is only matched once we add at least 2000 synthetic data points (an order or magnitude larger). On FairyTaleQA, we get enormous estimates for the number of additional synthetic points needed (a mean of 2.8e5). We do not interpret these numbers literally, rather seeing this as a sign that human generated data may occasionally boost performance to an extent that could be fundamentally unachievable by purely synthetic data.
@@ Appendix B Synthetic Data Generation
For the synthetic datasets used in our experiments, we correct for this label shift by specifying the label of the claim we wish to generate and providing only examples of claims with that specific label in the prompt. his setting is generous towards synthetic data generation. In practice, we might only have a fixed three examples to use in the prompt, potentially reducing the diversity of synthetic data generated. We may also not know the correct label proportion to ask for, and suffer a significant label shift when using synthetic data generation.
@@ Appendix C Datasets Used
All datasets used below are released under open use licenses, authorizing their use in this research. For each dataset below, we discuss the potential of dataset leakage (i.e., whether the data has been exposed to GPT-3.5-Turbo during its training) as well as the extent of automation involved in the generation of each dataset. However, across all experiments and ablations, these factors do not seem to have any discernable effect on the trends discovered in this work.
@@ C.1 Fact Verification Datasets
SciFact (Wadden et al., 2020) is a fact verification dataset for the scientific domain, which uses the abstracts of scientific articles as evidence texts. The corpus is collected from S2ORC (Lo et al., 2019), a publicly-available corpus of millions of scientific articles. Annotators are shown a source citation in the context of an article, and are asked to write up to three claims based on the content of the citation.
The above datasets are popular NLP challenge sets that were well known even before the release of the GPT-3.5-Turbo (Brown, 2020), the prompting model used in this work. The following two datasets were released after the official training date cut-off, guaranteeing that the data has not been seen ahead of time.
WANLI (Liu et al., 2022) is an NLI dataset of 108K examples created through a hybrid worker and AI collaboration approach. The creators first study MultiNLI (Williams et al., 2018) and use dataset cartography to automatically identify examples that demonstrate challenging reasoning patterns. and then instruct GPT-3 to compose new examples with similar patterns. Machine generated examples are then automatically filtered, and finally revised and labeled by human crowd workers. While GPT-3.5-Turbo has not been trained on this data, it is worth noting that the data is partially synthetic generated.
FACTIFY (Mishra et al., 2022) is a dataset on multi-modal fact verification. It contains images, textual claims, reference textual documents and reference images. The dataset marks some examples that can be verified using text only; we use this sample in our experiments. This dataset was released after the training cut-off date for GPT-3.5 and takes its evidence texts / claims from human-written news or editorial articles. This ensures that the prompt model studied have not seen the data before training.
@@ C.2 Question Answering Datasets
QAConv (Wu et al., 2021) focuses on informative conversations, including business emails, panel discussions, and work channels. The creators collect QA pairs with both human-written and machine-generated questions. They use a question generator and a dialogue summarizer as auxiliary tools to collect and recommend questions. While the arXiv version of the paper appears before the GPT-3 cut-off data (April 2021 to the cut-off date of Sept 2021), the paper itself appeared only at ACL2022. It is still possible that the training data was compromised, and owing to the lack of clarity on the training data used for GPT-3 we have no way to confirm or deny this speculation.
@@ Appendix D Detailed Discussion on Differences Between Synthetic and Human Data
To compute the extent to which the evidence sentences ‘contains’ the questions, answers, and claims, we measured the BLEU of the generation with each individual sentence of the evidence texts, plotting the maximum of these BLEU scores in Figure 17. We find that synthetic generations have a far higher n-gram overlap with the evidence sentences than human generations. This suggests that synthetic data generation produces data points that are more extractive, while humans are more likely to abstract from the evidence. We also use the position of the evidence sentence that achieves the highest BLEU score as a proxy for the source location of the synthetic generation, and find that synthetic data generation chooses more diverse sources for the question and answer content, with human annotation overwhelmingly more likely to create questions whose answers lie in the start of the evidence texts (Figure 18). Finally, the main text shows the size length comparison for a single dataset, here we provide a larger sample (Figure 16). We explored the errors created by the models trained on 0% and 100% data, searching for trends or divergences between the input instances that achieve a low prediction accuracy or score. Our investigation found no major distinguishing factors between them, leaving a more fine-grained study of the effect of purely synthetic data on model decision-making to future work.
@@ Appendix E Implementation Details
Hardware and Systems Used: The experiments were run on a cluster that included nodes with: 5 A40 GPUs (48GB), 3 RTX 2080Tis, and a separate machine using a single A100 GPU.
@@
Fine-Tuning Models Used: We used 2 fine-tuning models in our experiments. Llama3 used the Llama3.1-8B HuggingFace Checkpoint, and Mistral used the Mistral7B-Instruct-v0.2 HuggingFace Checkpoint. We did not conduct an extensive hyperparameter search, however we tried various epochs on smaller samples of the FEVER and ROPES datasets, selecting that number for every dataset on all experiments. Fact verification models used Adam Optimization with a learning rate of 1e-5 for 2 epochs, while QA datasets used a learning rate of 1e-2 for 5 epochs.
@@
Dataset Synthetic Fraction EM Inc R Inc BLEU ROUGE BERTScore
CoQA 0 40.6 52.2 60..8 47.85 64.01 87.78
0.25 35 51.8 54.8 44.05 61.95 85.89
0.5 31.6 42.8 60.6 38.68 54.65 79.88
0.75 39.2 50.4 69 46.38 62.07 79.94
0.9 36.2 50.2 58.2 44.46 60.81 85.4
1 13.6 26 58.2 18.17 26.55 52.76
FairytaleQA 0 0 0 0 39.34 55.3 90.82
0.25 0 0 0 40.02 56.12 90.5
0.5 0 0 0 39.22 55.3 88.9
0.75 0 0 0 39.74 55.36 90.56
0.9 0 0 0 38.1 54.2 89.9
1 0 0 0 30.12 49.51 88.35
QAConv 0 29.39 36.3 49.42 35.11 51.62 89.89
0.25 28.55 34.5 47.96 33.51 48.56 89.51
0.5 27.95 34.04 47.1 33.1 48.67 89.46
0.75 28.55 35.87 48.45 34.17 50.52 89.97
0.9 29.01 35.96 49.88 34.75 50.57 88.97
1 23.5 34.29 40.98 30.21 45.16 87.15
ROPES 0 66.76 67.35 72.15 67.01 72.66 95.95
0.25 66.82 67.53 69.78 67.15 71.12 96.15
0.5 62.79 63.44 65.46 63.07 66.61 95.17
0.75 66.82 68 68.83 67.39 70.82 96.2
0.9 70.57 71.46 71.82 70.98 73.01 96.84
1 60.84 63.86 61.19 62.07 64.9 95.29
@@
Run Dataset Synthetic Fraction BLEU ROUGE BERTScore
0 FairytaleQA 0.95 38.78 54.8 90.34
@@
Dataset Mean Median 25th Percentile 75th Percentile
WANLI 17671.7 16905.38 9711.42 22931.17
ROPES 17332.69 6006.45 3622.95 21944.36
FairyTaleQA 281951.01 36901.45 15128.74 813134.81
FEVER 1155.13 236.37 -1400.05 7073.19
Table 3: Summary statistics for the number of additional synthetic data points needed to match performance gain of 200 human data points (aggregated over various fixed operating points of synthetic data). High values for FairyTaleQA suggests that human generated data may unlock performance that is unachievable with purely synthetic data. Negative values for FEVER are due to a saturation of the performance gains, however human data helps reach the saturation point much faster (see Section A and Figure 14).

### v2

@@ 9 Ethical Considerations
The usage of synthetic data has several important ethical considerations. In the era of LMs trained on internet-wide corpora having poor documentation as to their exact data sources, it becomes challenging to ensure the privacy of individuals whose data may be obtainable via a public crawl (Yao et al., 2024). Additionally, models trained on massive internet-based data sources may contain implicit biases, as well as illegal and/or highly offensive material that is hard to audit and clean (Bender et al., 2021). This data affects the synthetic data obtained from prompt models, and could unknowingly impose cultural or ethical viewpoints that are unintended or not well aligned with the use case in mind. Specifically, prior work has shown that one of the prompt models studied in this work, GPT-3.5, often disagrees with humans on key ethical questions (Felkner et al., 2024). The endeavour to completely replace human annotation with synthetic data generation also has key implications on the extent to which the field of NLP employs human annotators. It is possible that an increasing reliance on purely synthetic data reduces the demand for human annotation, which would place a downward pressure on the working standards and compensation awarded to the remaining human annotators (Weidinger et al., 2022). We argue in this work that we should not try to eliminate human annotation from our dataset and method design, showing that their work contributes uniquely helpful data points.
@@ 10 Acknowledgements
This work was funded by the Defense Advanced Research Projects Agency with award HR00112220046. Any opinions, findings, conclusions, or recommendations expressed here are those of the authors and do not necessarily reflect the view of our sponsors.
@@ Appendix A Supplemental Figures
We present a detailed set of figures and tables to supplement the results presented in the main text.
@@
Main Experiments: For figures in the main text where only one task is shown (Figure 1 and Figure 2), we provide the complete figures with both tasks (Figure 6 and Figure 7). We also provide the individual performance curves for these experiments (Figure 8 and Figure 10).
Robustness to choice of QA metric: To verify the robustness of the results, we show that the QA results are not an artifact of the choice of metric (Figure 3 and Figure 4) by using Exact Match, String Inclusion, ROUGE-1 (Lin, 2004) and BERTScore (Zhang et al., 2020). There is overwhelming agreement between all metrics on the rankings of models.
Addressing spurious correlations: We show that the performance gains afforded by human generated data cannot be explained by a spurious correlation between the human generated train and test splits. This would occur when there are significant annotation artifacts that are not relevant to the task, but are correlated with the correct output. We conduct an out-of-domain experiment (Table 5), using different datasets to source the training data and testing on a single hold out dataset. Using more synthetic data leads to performance declines even in the OOD setting, showing that human data is of higher quality and the results from the main text cannot be explained by a spurious correlation between the human test and human training samples. Interestingly, in the OOD setting the decline is steady, and we do not observe the phenomenon of a small amount of human data having a disproportionate impact on performance. This suggests that the disproportionate impact of human data occurs when the human data is in-domain. We leave a further exploration of the OOD generalization abilities of synthetic vs. human data to future work.
Multilingual Experiments: We replicate our experiment using the Arabic, Georgian, and Indonesian splits of the XFact dataset. We observe (Figure 9) the same trend as those from earlier experiments, confirming that our results are not limited to the English language. While the phenomenon is reproduced, the threshold of replacement at which we observe a precipitous decline is not the same across languages. We hypothesize that the language-specific threshold at which a little human data leads to significant performance increases is dependent on how low resource the language is. The study of synthetic data in the multilingual setting has unique considerations that we have not addressed in this work; we leave a focus on these problems to future work.
@@
Ablations: We show that the same trends can be seen (Figures 11, 12, 13, 14 and 15) when using a different fine-tuning model (Mistral-7B), models of varying scales (from 1B parameter models to 30B parameter models. different prompting models (GPT-4 and Claude-3.5) and a more sophisticated prompting strategy (Chain-of-Thought Prompting). Across all configurations, we see a consistent decrease in performance when moving from 95% to 100% synthetic data, confirming that models trained on purely synthetic data can be improved by including just 125 real data points. For Chain-of-Thought Prompting, the authors manually annotated 3 examples with rationales per dataset to serve as the prompts. The complete examples and pipeline are provided with the code: github.com/dhananjayashok/littlehumandata
@@
We additionally show that these trends hold across data scales (Figures 16 and 17), replicating the experiment with n=3000 and n=1000. While the trend is clearly visible in both cases, the results for n=1000 have more variance and hence have a minority of cases where the relationship does not hold.
@@
Tradeoff Experiment: The main text shows results for the experiment detailed in Section 4 on the WANLI dataset (Figure 4), here we show results on the remaining three datasets (Figure 18) and provide (Figure 1) the number of additional synthetic points needed to match the performance gains of 200 additional real points (average, median and standard deviation for each dataset). ROPES shows similar results to WANLI, however FairyTaleQA and FEVER present different trends. On FEVER, we are able to reach the saturation point, after which additional data (whether synthetic or real) does not increase performance. Even in this case, we are able to reach this point of diminishing marginal return more rapidly when using a small amount of synthetic data. On a base synthetic training set of size $3000$, adding $200$ real data points drives the test accuracy to $89.25\%$, a score that is only matched once we add at least 2000 synthetic data points (an order of magnitude larger). On FairyTaleQA, we get enormous estimates for the number of additional synthetic points needed (a mean of 2.8e5). We do not interpret these numbers literally, rather seeing this as a sign that human generated data may occasionally boost performance to an extent that could be fundamentally unachievable by purely synthetic data.
@@ Appendix B Synthetic Data Generation
For the synthetic datasets used in our experiments, we correct for this label shift by specifying the label of the claim we wish to generate and providing only examples of claims with that specific label in the prompt.
For all datasets, we verify that the diversity of the generated claims/questions/answers are comparable to that of the human generated texts (see Appendix D)
This setting is generous towards synthetic data generation. In practice, we might only have three fixed examples to use in the prompt, potentially reducing the diversity of synthetic data generated. We verify that this does not affect the results of our experiment in the Chain-of-Thought ablation (Figure 15), where we use a fixed set of examples to generate all synthetic data points. We may also not know the correct label proportion to ask for and suffer a significant label shift when using synthetic data generation.
@@ Appendix C Datasets Used
All datasets used below are released under open use licenses, authorizing their use in this research. For each dataset, we discuss the potential of dataset leakage (i.e., whether the data has been exposed to GPT-3.5-Turbo during its training) as well as the extent of automation involved in the generation of each dataset. However, across all experiments and ablations, these factors do not seem to have any discernable effect on the trends discovered in this work.
@@ C.1 Fact Verification Datasets
SciFact (Wadden et al., 2020) is a fact verification dataset for the scientific domain, which uses the abstracts of scientific articles as evidence texts. The corpus is collected from S2ORC (Lo et al., 2020), a publicly-available corpus of millions of scientific articles. Annotators are shown a source citation in the context of an article, and are asked to write up to three claims based on the content of the citation.
The above datasets are popular NLP challenge sets that were well known even before the release of GPT-3.5-Turbo (Brown et al., 2020), the prompting model used in this work. The following two datasets were released after the official training date cut-off, guaranteeing that the data has not been seen ahead of time.
WANLI (Liu et al., 2022) is an NLI dataset of 108K examples created through a hybrid worker and AI collaboration approach. The creators first study MultiNLI (Williams et al., 2018) and use dataset cartography to automatically identify examples that demonstrate challenging reasoning patterns. and then instruct GPT-3 to compose new examples with similar patterns. Machine generated examples are then automatically filtered, and finally revised and labeled by human crowd workers. While GPT-3.5-Turbo has not been trained on this data, it is worth noting that the data is partially synthetically generated.
FACTIFY (Mishra et al., 2022) is a dataset on multi-modal fact verification. It contains images, textual claims, reference textual documents and reference images. The dataset marks some examples that can be verified using text only; we use this sample in our experiments. This dataset was released after the training cut-off date for GPT-3.5 and takes its evidence textsclaims from human-written news or editorial articles. This ensures that the prompt models studied have not seen the data before training.
@@ C.2 Question Answering Datasets
QAConv (Wu et al., 2022) focuses on informative conversations, including business emails, panel discussions, and work channels. The creators collect QA pairs with both human-written and machine-generated questions. They use a question generator and a dialogue summarizer as auxiliary tools to collect and recommend questions. While the arXiv version of the paper appears before the GPT-3 cut-off data (April 2021 to the cut-off date of Sept 2021), the paper itself appeared only at ACL 2022. It is still possible that the training data was compromised, and owing to the lack of clarity on the training data used for GPT-3 we have no way to confirm or deny this speculation.
@@ Appendix D Detailed Discussion on Differences Between Synthetic and Human Data
To compute the extent to which the evidence sentences ‘contain’ the questions, answers, and claims, we measure the BLEU of the generation with each individual sentence of the evidence texts, plotting the maximum of these BLEU scores in Figure 21. We find that synthetic generations have a far higher n-gram overlap with the evidence sentences than human generations. This suggests that synthetic data generation produces data points that are more extractive, while humans are more likely to abstract from the evidence. We also use the position of the evidence sentence that achieves the highest BLEU score as a proxy for the source location of the synthetic generation, and find that synthetic data generation chooses more diverse sources for the question and answer content, with human annotation overwhelmingly more likely to create questions whose answers lie in the start of the evidence texts (Figure 22). Finally, the main text shows the size length comparison for a single dataset. Here we provide a larger sample (Figure 20). We explore the errors created by the models trained on 0% and 100% data, searching for trends or divergences between the input instances that achieve a low prediction accuracy or score. Our investigation finds no major distinguishing factors between them, leaving a more fine-grained study of the effect of purely synthetic data on model decision-making to future work.
@@ Appendix E Implementation Details
Hardware and Systems Used: The experiments were run on a cluster that included nodes with: five A40 GPUs (48GB), three RTX 2080Tis, and a separate machine using a single A100 GPU.
@@
Fine-Tuning Models Used: We used two fine-tuning models in our experiments. Llama3 used the Llama3.1-8B HuggingFace Checkpoint, and Mistral used the Mistral7B-Instruct-v0.2 HuggingFace Checkpoint. We did not conduct an extensive hyperparameter search, however we tried various epochs on smaller samples of the FEVER and ROPES datasets, selecting that number for every dataset on all experiments. Fact verification models used Adam Optimization with a learning rate of 1e-5 for two epochs, while QA datasets used a learning rate of 1e-2 for five epochs.
@@
Dataset Synthetic % EM Inc R Inc BLEU ROUGE BERTScore
CoQA 0 40.6 52.2 60.8 47.9 64.0 87.8
25 35.0 51.8 54.8 44.1 62.0 85.9
50 31.6 42.8 60.6 38.7 54.7 79.9
75 39.2 50.4 69.0 46.4 62.1 79.9
90 36.2 50.2 58.2 44.5 60.8 85.4
100 13.6 26.0 58.2 18.2 26.6 52.8
FairytaleQA 0 0.0 0.0 0.0 39.3 55.3 90.8
25 0.0 0.0 0.0 40.0 56.1 90.5
50 0.0 0.0 0.0 39.2 55.3 88.9
75 0.0 0.0 0.0 39.7 55.4 90.6
90 0.0 0.0 0.0 38.1 54.2 89.9
100 0.0 0.0 0.0 30.1 49.5 88.4
QAConv 0 29.4 36.3 49.4 35.1 51.6 89.9
25 28.6 34.5 48.0 33.5 48.7 89.6
50 28.0 34.0 47.1 33.1 48.7 89.5
75 28.6 35.9 48.5 34.2 50.5 90.0
90 29.0 36.0 49.9 34.8 50.6 89.0
100 23.5 34.3 41.0 30.2 45.2 87.2
ROPES 0 66.8 67.4 72.2 67.0 72.7 96.0
25 66.8 67.5 69.8 67.2 71.1 96.2
50 62.8 63.4 65.5 63.1 66.6 95.2
75 66.8 68.0 68.8 67.4 70.8 96.2
90 70.6 71.5 71.8 71.0 73.0 96.8
100 60.8 63.9 61.2 62.1 64.9 95.3
@@
Run Dataset Synthetic % BLEU ROUGE BERTScore
0 FairytaleQA 95 38.78 54.80 90.34
@@
Train Sets Test Set Synthetic % Test Accuracy
FEVER, SciFact WANLI 0 69.98
25 67.56
50 65.86
75 64.82
100 64.36
WANLI, SciFact FEVER 0 83.01
25 80.64
50 79.22
75 78.94
100 76.01
FEVER, WANLI SciFact 0 71.76
25 69.75
50 69.82
75 66.42
100 64.41
Table 5: Test accuracy when replacing human data with synthetic data in the out-of-distribution setting. Using more synthetic data leads to performance declines even in the OOD setting, showing that human data is of higher quality and the results from the main text cannot be explained by a spurious correlation between the human test and human training samples.
Figure 9: Change in model performance as the proportion of synthetic points in the training data is increased on multilingual fact verification datasets (splits of X-Fact). We observe the same trend as those from earlier experiments, confirming that our results are not limited to the English language. While the phenomenon is reproduced, the threshold of replacement at which we observe a precipitous decline is not the same across languages.
@@
Figure 12: Results hold when using Claude-3.5-Sonnet as the prompting model, showing that the phenomenon is not particular to Synthetic Data from GPT based models.
Figure 13: Results hold consistently on Fact Verification datasets when using models of different scales.
@@
Dataset Claim / Question
Synthetic Human
FEVER 35.78 42.76
WANLI 15.15 20.10
SCIFACT 7.12 20.92
FACTIFY 14.50 23.93
NarrativeQA 30.92 8.25
CoQA 7.08 8.39
FairyTaleQA 22.59 16.85
ROPES 28.73 41.42
Table 6: 4-Gram overlap % between all synthetic and human generated claims / questions for each dataset. On several datasets, synthetic claims have a lower overlap